### v2

# HarmonRank: Ranking-aligned Multi-objective Ensemble for Live-streaming E-commerce Recommendation

Abstract.
Recommendation for live-streaming e-commerce is gaining increasing attention due to the explosive growth of the live streaming economy. Different from traditional e-commerce, live-streaming e-commerce shifts the focus from products to streamers. This shift requires the ranking mechanism to balance both purchases (short-term value) and user-streamer interactions (long-term value), e.g., follows and comments, for the ecology health of the platform. To trade off multiple objectives, a popular solution is to build an ensemble model to integrate multi-objective scores into a unified score. The ensemble model is usually supervised by multiple independent binary classification losses of all objectives. However, this paradigm suffers from two inherent limitations. First, the optimization direction of the binary classification task is misaligned with the ranking task (evaluated by AUC). Second, this paradigm overlooks the alignment between objectives, e.g., comment and buy behaviors are partially dependent which can be revealed in labels correlations. The model can achieve better trade-offs if it learns the aligned parts of ranking abilities among different objectives.
To mitigate these limitations, we propose a novel multi-objective ensemble framework HarmonRank to fulfill both alignment to the ranking task and alignment among objectives. For alignment to ranking, we formulate ranking metric AUC as a rank-sum problem and utilize differentiable ranking techniques for ranking-oriented optimization. For inter-objective alignment, we change the original one-step ensemble paradigm to a two-step one. In 1st step, we align each objective with other objectives in a self-attention mechanism. In 2nd step, we fuse aligned objective encodings into the ensemble score with the guidance of personalized features.
Extensive offline experiments results on two industrial datasets and online experiments demonstrate that our approach significantly outperforms existing state-of-the-art methods. The proposed method has been fully deployed in Kuaishou’s live-streaming e-commerce recommendation platform with 400 million DAUs, contributing over 2% purchase gain.
Multi-objective ranking ensemble, Personalized recommendation, Learning to rank
$c$s: Information systems Recommender systems.
1. Introduction
Figure 1. Conceptual view of the multi-objective ensemble stage and different ensemble paradigms. (a) The position of multi-objective ensemble stage in industrial recommender systems. (b) Existing one-step ensemble paradigm based on multi-layer non-linear with multiple binary classification losses. (c) Limitations of existing paradigm. First, lack of the alignment to ranking task. Second, lack of the alignment between objectives.(d) Our proposed paradigm to fulfill alignment to ranking and inter-objective alignments.
Live-streaming e-commerce market has recently experienced explosive growth. Different from traditional e-commerce, the products are endorsed and promoted by streamers in live-streaming e-commerce, which shifts the focus from products to the streamer. This paradigm shift necessitates a change from purchase-oriented ranking mechanism to multi-objective balanced mechanism including interactions between users and streamers, e.g., follows and comments. Different from short-term purchase, these interaction-oriented objectives matter to the ecosystem for the their long-term value. For an example, follow behaviors can foster long-term user-streamer relationships, which helps to the ecology health of the platform. This change poses more challenges for balancing these competing objectives than traditional purchase-driven e-commerce recommender systems.
To balance these competing objectives, industrial ranking systems usually follow a two-stage paradigm (see Fig. 1 (a)), i.e., multi-objective prediction (MP) and multi-objective ensemble (ME) (Pei et al., 2019; Cao et al., 2025; Li et al., 2023; Carmel et al., 2020). The MP stage is usually a high-capacity multi-task learning model (Ma et al., 2018a) whose goal is to predict the precise probability scores for multiple objectives, e.g., click through rate (CTR). Whereas ME stage is a light-weight model whose goal is to generate a unified ensemble score in a personalized manner. It usually only takes aforementioned multi-objective scores and a few user features as inputs. The ensemble score is used to truncate top items to represent for users. In this paper, we focus on ME problem. Similar to MP stage, in ME stage we use AUC (Area Under ROC Curve) (Bamber, 1975) between the ensemble score and the ground-truth labels to evaluate the ranking abilities of the ensemble model. The only difference is that we utilize the sum of AUC of all objectives as the metic for comprehensive multi-objective ranking evaluation.
The key challenge at this stage lies in the absence of a gold-standard supervision to comprehensively assess items considering multiple objectives. To mitigate this problem, a typical solution is to impose multiple independent binary classification supervisions over the ensemble score (Cao et al., 2025; Lin et al., 2019) (see Fig. 1 (b)). However, this paradigm suffers from two limitations, for which we present a conceptual view in Fig. 1 (c). (1) Lack of alignment to ranking task. The optimization directions of binary classification based supervisions are misaligned with the ultimate ranking task (evaluated by AUC) (Tang et al., 2022). For an example, if the positive labels of some action are highly sparse, a dummy predictor which predicts all instances as negative seems to present an acceptable binary classification result. But it actually yields a poor ranking results with an AUC of 0.5. (2) Lack of inter-objective alignment. The different objectives are neither independent nor strictly adversarial - they are partially dependent. For an example, buy actions empirically exhibit high correlations with comment actions, which can be verified through the statistical analysis of correlation between objectives (see Fig.4). Thus, if model can learn the aligned parts among multi-objectives, it is easier for model to trade-off these competing objectives.
To mitigate these issues, in this paper, we propose a novel personalized multi-objective ensemble framework HarmonRank, which enables both alignment to the ranking task and alignment between multiple objectives. For alignment to ranking, we firstly reformulate the non-differentiable AUC as a rank sum problem. Then we utilize an advanced differentiable ranking technique to optimize multi-objective AUC in an end-to-end manner. For alignment between objectives, we propose to change existing one-step ensemble architecture to a two-step one, i.e., first align then ensemble. In align step, we align the common parts between multiple objectives via self-attention mechanism. In ensemble step, we fuse the aligned encodings of all objectives into ensemble score. In this process, we compress personalized information into a query and enable user-specific ensemble by cross attention mechanism.
Overall, our contributions can be summarized as follows:
  • We propose a novel personalized multi-objective ensemble framework HarmonRank to enable the alignment of optimization to the ranking task with differentiable AUC optimization.
  • We propose a two-step ensemble paradigm to enable align the shared parts between multiple objectives, to obtain better trade-off between objectives.
  • We conduct extensive offline experiments on two industrial recommendation datasets. Experimental results demonstrate that our HarmonRank achieves superior performance over all existing state-of-the-art methods. Additionally, HarmonRank contributes more than 2% purchase gain for Kuaishou live-streaming e-commerce platform.
2. Related Work
2.1. Multi-objective optimization in recommender systems
Multi-objective optimization is recognized as a key challenge in both research and industry community. Early recommender systems primarily optimized for only few objectives, e.g. click-through rate (CTR) prediction for e-commerce recommender system and watch time for video recommender systems. However, with the increasing complexity user interface functionalities, morden recommender system need to balance more competing objectives. Some studies attend on the sample selection bias caused by chain relationships between objectives, e.g., CTR and conversion rate (CVR) (Ma et al., 2018b; Cheng et al., 2025). Some other researchers are dedicated to improve multi-task learning mechanism between different objectives (Ma et al., 2018a; Tang et al., 2020). For an example, Multi-gate Mixture-of-Experts (MMOE) (Ma et al., 2018a) architecture addresses this by leveraging expert networks to capture shared representations across objectives, thereby improving accuracy of individual objectives. Despite these advances, how to ensemble multiple prediction scores of different objectives remains a non-trivial challenge. Current industrial systems typically address this through a multi-objective ensemble model (Li et al., 2023; Yu et al., 2024; Pei et al., 2019).
2.2. Multi-objective Ensemble in recommender systems
Early industrial approaches predominantly employed ranking formulas to combine scores from multiple objectives, typically optimizing parameters through grid search. While black-box optimization methods, e.g., Bayesian optimization (Močkus, 1974) and cross entropy methods (Rubinstein and Kroese, 2004), can reduce the iterations for searching optimum, they still suffer from two critical limitations: (1) inability to present personalized ranking mechanism to adapt user preferences, and (2) incapacity to exploit real-time user feedback. The factors matter for system since the optimal ensemble schemes change with different users and different time. To address these challenges, recent methods have proposed model-based ensemble solutions, to utilize more informative features and deep neural networks trained in hourly updated streaming feedback.
The key challenge for model-based multi-objective ensemble lies in the absence of a gold-standard supervision across multiple objectives. To solve this problem, some methods apply multiple unilateral binary classification losses to optimize towards all objectives simultaneously (Cao et al., 2025; Lin et al., 2019), which can be referred as loss aggregation methods. The other methods aim to build a unified supervision by aggregating multiple binary labels into a single regression label (Li et al., 2023; Carmel et al., 2020), which can be categorized as label aggregation methods. Despite these efforts, the optimization directions of these methods’ supervisions are misaligned with the ultimate ranking task (evaluated by AUC), which leads to suboptimal performance. Reinforce learning based methods (Pei et al., 2019) can directly optimize towards ranking task metrics (AUC sum), whereas their policy gradients suffer from large variance and instability. Few of existing methods consider to optimize final sum of AUC in an end-to-end differentiable manner. It is NP-hard to direct optimize AUC because of the non-convexity and discontinuousness of (Gao and Zhou, 2012).
2.3. AUC optimization
The AUC is a ranking quality metric defined based on pairwise comparisons (Bamber, 1975; Hanley and McNeil, 1982a) under binary label settings. Thus, traditional differentiable AUC optimization methods concentrate on developing effective pairwise surrogate functions to approximate non-differentiable comparison function, i.e., $s_{i}-s_{j}\geq 0$. For examples, pairwise square methods (Gao et al., 2013) employ square functions $(1-(s_{i}-s_{j}))^{2}$ surrogate functions, while pairwise logistic methods (Gao and Zhou, 2012) propose to use logistic function $\text{log}(1+\text{exp}(-(s_{i}-s_{j})))$ to approximate the comparison function. However, the inherent $O(n^{2})$ time complexity of pairwise approaches renders them impractical for large-scale datasets. Consequently, numerous optimization attempts have shifted towards instance-wise AUC optimization. Leveraging the special form of pairwise square, pairwise AUC maximization can be decomposed into a min-max instance-wise formulation (Ying et al., 2016), with $O(N)$ time complexity. Nevertheless, this instance-wise optimization introduces much sensitivity to noisy samples. To address this, AUCM (Yuan et al., 2021) proposed a margin-based min-max optimization strategy to effectively mitigates this issue.
While recent advancements improve the computational complexity to $O(N)$, they inadvertently introduce greater inconsistency with factual AUC computation due to computation simplification for surrogate functions. In this paper, to balance computational efficiency and AUC computation consistency, we propose a novel AUC optimization method under a rank based formulation other than pairwise formulation. We formulate AUC computation as a rank-sum problem in the ordered score list and utlize differenable ranking technique to solve this rank-sum problem. Our method presents $O(n\text{log}n)$ time complexity and superior AUC optimization accuracy for alignment with AUC theoretical computation.
3. Problem Formulation
Figure 2. Systematic overview of our HarmonRank framewotk. (a) The architecture of HarmonRank. (b) Relation-aware module with inter-objective alignment and personalized guidance. (c) Relation-agnostic module with gate mechanism and linear fusion pathway. ‘pBuy’, ‘pCmt’, ‘pFollow’, ‘pLV’ and ‘pLike’ here represent the predicted probability scores corresponding to ‘buy’, ‘comment’, ‘follow’, ‘long view’ and ‘like’.
3.1. Input and output
Let $\mathcal{I}_{cand}(\mathbf{u},\mathbf{c})=\left\{I_{i}\right\}_{i}^{N}$ be a requested candidate set for the user $\mathbf{u}$ and environment context $\mathbf{c}$ (e.g., time and app version). For each item $I_{i}$, multi-objective prediction (MP) model estimates $M$ kind of interaction scores $\mathbf{e}_{i}\in\mathbb{R}^{[M,1]}$, e.g., purchase, long view, like, follow, etc. Multi-objective ensemble (ME) model is responsible for integrating these scores into an ensemble score in a personalized way:
(1) $s_{i}=\mathcal{F}(\mathbf{e}_{i},\mathbf{u},\mathbf{c};\theta)$
Then top items are selected and displayed to users based on top-k truncation of the ensemble scores $\mathcal{S}=\{s_{i}\}_{i}^{N}$, where $\mathcal{F}(\cdot;\theta)$ is a neural-network ensemble model parameterized by $\theta$.
(2) $\mathcal{I}_{realshow}=\text{TopK}(\mathcal{I}_{cand},\mathcal{S})$
3.2. Train and evaluation
Similar to MP stage, we collect multiple behavior feedback labels in a fixed waiting window for each exposed sample in logs, and we use these labels to supervise the ensemble model training. Our goal is to achieve improved trade-offs across competing objectives. Due to the high cost of online AB metric evaluation, we typically evaluate the ME model’s comprehensive ranking abilities on all objectives by the sum of AUC on a offline dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$. Here $\mathbf{y}_{i}\in\{0,1\}^{[M,1]}$ is the ground truth interaction labels for the $i$-th sample:
(3) $\mathcal{G}_{m}(\mathcal{D})=\frac{\sum_{s_{i}\in\mathcal{D}_{m}^{+},s_{j}\in\mathcal{D}_{m}^{-}}\mathbb{I}(s_{i}\geq s_{j})}{|\mathcal{D}_{m}^{+}|\cdot|\mathcal{D}_{m}^{-}|}$
(4) $\mathcal{G}(\mathcal{D})=\sum_{m=1}^{M}\mathcal{G}_{m}(\mathcal{D})$
where $\mathcal{D}_{m}^{+}$ and $\mathcal{D}_{m}^{-}$ represent the positive and negative sample sets defined on the $m$-th objective, e.g., the purchased items and unpurchased items. Existing methods primarily rely on classification or regression losses to supervise ME models, yet their optimization directions are substantial misaligned with the factual ranking ability (AUC). Therefore we propose an end-to-end ranking driven ME paradigm that explicitly optimizes the multi-objective AUC metric:
(5) $\mathbf{\mathbf{\theta}}^{*}=\text{argmax}_{\mathbf{\mathbf{\theta}}}\mathcal{G}(\mathbf{\mathbf{\theta}})$
4. HarmonRank
4.1. Alignment to ranking
4.1.1. Rank-based AUC formulation
Since $\mathcal{G}$ is non-differentiable w.r.t. $\mathbf{\theta}$, existing wisdom usually maximize AUC via surrogate function to approximate pairwise comparing function $\mathbb{I}(s_{i}\geq s_{j})$ i.e., pairwise logistic (Gao and Zhou, 2012) and pairwise square (Gao et al., 2013) functions. However, this formulation brings large inconsistency between physical AUC computation. Meanwhile, pairwise computation introduce quadratic computation cost during training, which hinders further scaling on datasets. To avoid these problems, we resort to alternative formulation for AUC computation.
In this section, we temporarily omit the subscripts $m$ of objectives for simplicity. When sorting all positive and negative samples in an increasing order, AUC can be simplified to compute the rank sum of positive samples for the good property of ordered list (Hanley and McNeil, 1982b):
(6) $\mathcal{G}(\mathbf{w})=\frac{\langle\mathbf{r},\mathbf{y}\rangle-|\mathcal{D}^{+}|\cdot(|\mathcal{D}^{+}|+1)/2}{|\mathcal{D}^{+}|\cdot|\mathcal{D}^{-}|}$
where the vector $\mathbf{r}=\text{Rank}(\mathbf{s})$ represents the ranks of all $N$ samples and $\mathbf{y}$. It is noteworthy that, the conventional pairwise formulation suffers from cost inefficiency because of additional computations for positive-negative pairs with reverse orders ($s_{+}<s_{-}$). Under the ordered setting, AUC computation can be simplified to $O(nlogn)$ time complexity. To optimize $\mathcal{G}(\mathbf{w})$, we only need to optimize $\langle\mathbf{r},\mathbf{y}\rangle$ for other terms are constant numbers. So the objective can be simplified as:
(7) $\mathbf{\mathbf{\theta}}^{*}=\text{argmax}_{\mathbf{\mathbf{\theta}}}\langle\mathbf{r},\mathbf{y}\rangle$
However, for rank operation $\text{Rank}(\cdot)$ is non-convex, discontinuous gradients is infeasible for $\mathbf{s}$. To alleviate this, we employ a fast differentiable ranking algorithm (Blondel et al., 2020) to enable gradient backpropagation from supervisions to the ensemble model.
4.1.2. Differentiable Ranking
Following (Blondel et al., 2020), we cast ranking operation (in increasing order here) as a discrete optimization problem over all feasible permutations $\Sigma$.
(8) $\mathbf{r}^{*}=\text{argmax}_{\sigma\in\Sigma}\langle\mathbf{s},\mathbf{r_{\sigma}}\rangle$
where $|\Sigma|=A^{N}_{N}$ and we desire the earlier-ranked samples should have smaller values. To transform this discrete optimization form to a continuous optimization problem, this method introduces a convex hull composed of all feasible permutations in $\Sigma$, which forms a permutahedron $\mathcal{P}(\mathbf{r})$:
(9) $\mathcal{P}(\mathbf{r}):=\text{conv}({r_{\sigma}}|\sigma\in\Sigma)$
To make it a convex objective, this method employs a quadratic regularization into the original optimization problem:
(10) $P(\mathbf{s},\mathbf{r})=\text{argmax}_{\mathbf{t}\in\mathcal{P}(\mathbf{r})}\langle\mathbf{s},\mathbf{t}\rangle-\frac{1}{2}\parallel\mathbf{t}\parallel^{2}=\text{argmin}_{\mathbf{t}\in\mathcal{P}(\mathbf{r})}\frac{1}{2}\parallel\mathbf{t}-\mathbf{s}\parallel^{2}$
In this way, the original problem has been transformed to the projections to the permutahedron, which is a strong convex function. This approach enables forward propagation with $O(n\text{log}n)$ time complexity for sorting operation. In the same time, it cost only $O(n)$ time complexity during backward propagation for the gradients computation can be finished in the sorted sequence (Blondel et al., 2020).
4.2. Alignment among objectives
We find prior wisdom consistently overlook the alignment between multiple objectives, which makes model hard to learn the shared ranking abilities among objectives. As shown in Fig. 4, different objectives can be partially aligned with varying degrees. To fill this gap, we design a dual-path architecture to integrate multiple scores in both one-step (relation-agnostic) manner and two-step (relation-aware) manner between objectives. As shown in Fig. 2 (a), we split our framework into two parallel components: relation-aware module and relation-agnostic module. The output scores of two modules are lately fused into a united score to compute the AUC loss.
4.2.1. Relation-aware Module
We align the shared parts between objectives with a relation-aware two-step paradigm, first align then ensemble. For align step, as shown in Fig. 2(b), we utilize self-attention mechanism (Vaswani et al., 2017) to capture pairwise relations between objectives. Through dynamic computation of relations between objectives, the aligned common parts between them are extracted. Specifically, the softmax normalized dot products between projected objective vectors, $\mathbf{Q}_{r}$ and $\mathbf{K}_{r}$, form the self attention matrix $\mathbf{A}_{r}\in\mathbb{R}^{m\times m}$ between them. This matrix naturally represents the relation strengths between objectives. Then the original score embeddings are integrated into the representations $\mathbf{x}_{r}$ according to the attention matrix.
(11) $\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r}$ $=\mathbf{x}\mathbf{W}^{Q_{r}},\mathbf{x}\mathbf{W}^{K_{r}},\mathbf{x}\mathbf{W}^{V_{r}}$
(12) $\mathbf{A}_{r}$ $=\text{Softmax}(\mathbf{Q}_{s}\mathbf{K}_{s}^{\top}/\sqrt{d_{k}})$
(13) $\mathbf{x}_{r}$ $=\mathbf{A}_{r}\mathbf{V}_{r}$
Although we achieve inter-objective alignment, we still lack clear guidance on how to fuse multiple objectives into a unified score. Fortunately, user-specific personalized information, including user profile and context feature, can serve as strong inductive bias to guide this fusion. We cast the personalized information as a query, and search with it over multi-objective scores. We use the searched result as the fused score encodings to obtain ensemble score. In this way, we implicitly let model answer a question that ‘Which objects can best represent the user’s holistic intentions in the current context?’. Specifically, we transform personalized information into the query vector $\mathbf{Q}_{p}$ by linear projection, similar to key and value vectors.
(14) $\mathbf{Q}_{p},\mathbf{K}_{p},\mathbf{V}_{p}$ $=\mathbf{P}\mathbf{W}^{Q_{p}},\mathbf{x}_{r}\mathbf{W}^{K_{p}},\mathbf{x}_{r}\mathbf{W}^{V_{p}}$
(15) $\mathbf{A}_{p}$ $=\text{Softmax}(\mathbf{Q}_{p}\mathbf{K}_{p}^{\top}/\sqrt{d_{k}})$
Then attention weights $\mathbf{A}_{r}\in\mathbb{R}^{1\times m}$ , computed via dot product over all objective vectors $\mathbf{s}_{r}$, can serve as personalized guidance for importance of objectives for ensemble models. Then we fused score encodings by attention weights and transform the output to scalar ensemble score $\mathbf{s}_{1}\in\mathbb{R}$ by a linear projection layer.
(16) $\mathbf{s}_{1}=\mathbf{w}_{1}^{\top}\left(\mathbf{A}_{p}\mathbf{V}_{p}\right)+b_{1}$
4.2.2. Relation-agnostic Module
Although we have effectively align the shared parts between different objectives, we still need to ensure that original information of each objective score can be fully preserved in final ensemble process. To this end, we introduce a relation-agnostic module (see Fig. 2(c)). In this module, we employ a gate mechanism to dynamically adjust the importance of different objectives based on their input scores. We obtain the gating coefficients $\mathbf{g}\in[0,1]^{m}$ by projecting score encodings through a linear layer with sigmoid activations. We control the score-wise retaining ratio of all score encodings through Hadamard product between coefficients and score encodings. We then compute the output ensemble score by a final linear fusion layer over gated encodings.
(17) $\mathbf{g}$ $=\text{Sigmoid}(\mathbf{w}_{g}^{\top}\mathbf{x}+b_{g})$
(18) $\mathbf{s}_{2}$ $=\mathbf{w}^{\top}_{2}(\mathbf{g}\odot\mathbf{x})+b_{2}$
Simultaneously, to make ensemble model to learn a first-order linear fusion scheme, we maintain a parallel linear fusion pathway. In this way, we prevent model performance degradation when we add previous high-order compositions between objectives and make the whole ensemble model more robust.
(19) $\mathbf{s}_{3}$ $=\mathbf{w}_{3}^{\top}\mathbf{x}+b_{3}$
Finally, we obtain the output ensemble score by additive fusion of the outputs of both relation-aware module ($\mathbf{s}_{1}$) and relation agnostic module ($\mathbf{s}_{2}$ and $\mathbf{s}_{3}$).
(20) $\mathbf{s}$ $=\mathbf{s}_{1}+\mathbf{s}_{2}+\mathbf{s}_{3}$
4.3. Pre-processing
Before feed the score into relation-aware and relation-agnostic modules. We apply discretization embedding to represent the multi-objective scores to obtain non-linear representation. We use a simple equal distance discretization embedding technique to transform scalar scores to embeddings (Guo et al., 2021).
(21) $\mathbf{e}^{m}=\mathbf{E}^{m}\cdot\text{floor}(\frac{\mathbf{s}^{m}}{B})$
where $B$ is the number of discrete buckets and $\mathbf{E}^{m}$ is the embedding matrix for the $m$th objective. The final inputs for the ranking model is the concatenation of both these discrete embeddings: $\mathbf{x}=[\mathbf{e}^{1};...\mathbf{e}^{m}]$. In this way, we empower the model more non-linear fitting power comparing to simply using the scalar score representations.
5. Experiments
TenRec-QKVideo Kuaishou-ELive
method 3-objectives 5-objectives 3-objectives 5-objectives
Multi-objective BCE (Cao et al., 2025) 2.2420 3.6465 2.1555 3.6682
Label aggregation (Li et al., 2023) 2.4040 3.8311 2.0765 3.6883
Reinforce learning (Pei et al., 2019) 2.2488 3.9074 2.1930 3.7467
Ours 2.4617 3.9186 2.2191 3.7522
Table 1. Offline results on AUC sum on two industrial datasets under two settings. For TenRec-QKVideo, ‘3 objectives’ are click, like and follow, while ‘5 objectives’ include click, like, follow, share and long view. For Kuaishou-ELive, the ‘3 objectives’ are buy, follow and long view, while ‘5 objectives’ include buy, follow, like, comment and long view.
In this section, we conduct extensive online and offline experiments to verify the efficacy of our model following 4 research questions:
  • RQ1: How does our model perform on industrial recommenda- tion datasets?
  • RQ2: How our module design and hyperparameter choice impact the performance?
  • RQ3: Can our method bring improvements of A/B test metrics on online product environment?
  • RQ4: How our method impact trade-off between objectives? What do the model learn in multiple objective alignment?
5.1. Implementation Details
Before training ensemble model, we firstly prepare the input multi-objective scores via performing inference with a high-capacity ranking model over both datasets. For Kuaishou-Elive, we use the online in-service model trained incrementally by years of logs. For the public TenRec dataset, we train a standard MMOE (Ma et al., 2018a) ranking model. Based on prepared multi-objective scores, all offline models are trained with 500 epochs with SGD optimizer. In terms of online training model with streaming data, it is usually trained for one epoch for the consideration of run-time cost and one-epoch overfiting risk in ranking stage. This scheme results in low sample efficiency. Since the model of ME stage is much more lightweight than MP stage and no sparse ID feature is used (cause of one-epoch overfitting), we train 20 epochs for each batch of streaming samples, to balance run-time cost and sample efficiency. We set learning rate to 1e-4 and batch size = 10240 for all methods.
5.2. Datasets
To evaluate the performance of our method and compared baselines, we conduct comprehensive experiments on two real-world industrial datasets.
Since there are few public live-streaming e-commerce dataset, we build a new dataset called Kuaishou E-Live, which contains both e-commerce feedback (purchase) and engagement oriented feedbacks (long view, like, follow and comment). We built it by sampling millions of exposure logs at Kuaishou e-commerce platform.
Moreover, to verify the universality of over method, we conduct experiments on a public short video recommendation dataset, TecRec-QKVideo (Yuan et al., 2022). TecRec-QKVideo is a collected from QQ-KAN video feeds platforms. It contains over 1M user interaction logs including five type of feedbacks, click, like, share, follow and long view. For the need of personalized ensemble, we utilize several user profile features like age, gender, etc. We train our method and compared methods with the same equal loss weights (all weights set to 1.0) for fair comparison on objective trade-offs.
5.3. Compared Methods.
We compare our methods with most representative methods on multi-objective ensemble task. For fair comparison, we use the same network structure as HarmonRank on these baselines.
  • Multi-objective binary cross entropy based (M-BCE) (Cao et al., 2025): This method enables optimization towards multiple objectives by integrate joint learning under multiple independent binary cross entropy losses for different objectives.
  • Reinforce learning (RL) based (Pei et al., 2019) : RL-based method formulate multi-objective ensemble as a markov decision process, and treat input multi-objective scores as the ‘state’ and the multi-object fusion score as the ‘action’. The multi-object AUC over a batch of data serve as ‘reward’.
  • Label aggregation based (Carmel et al., 2020; Li et al., 2023) : Label aggregation based method transforms multi-objective learning into single-objective learning by fusion multiple binary labels to a unified regression label. Then it utilize mean square error loss between the ensemble score and the aggregated label to optimize the ensemble model.
5.4. Offline Results (RQ1)
We evaluate the efficacy of our model by the multi-objective AUC sum. We present the results on two settings, i.e., 3 objectives and all 5 objectives, to verify the conclusion consistency over different settings. For TenRec-QKVideo, ‘3 objectives’ are click, like and follow, while ‘5 objectives’ include click, like, follow, share and long view. For Kuaishou-ELive, the ‘3 objectives’ are buy, follow and long view, while ‘5 objectives’ include buy, follow, like, comment and long view.
The experimental results on two datasets are shown in Tab. 1. On the TenRec-QKVideo dataset, our proposed method outperforms MBCE (Cao et al., 2025) by 27.2 pp (3.919 v.s. 3.647) for 5-objective AUC sum. On the Kuaishou-Elive dataset, our model outperforms MBCE by 8.4 pp (3.7522 v.s. 3.6682) for 5-objective AUC sum. The comparison results to the MBCE model demonstrates the efficacy of our proposed ranking-aligned ensemble model.
The RL-based model is the best performing baseline on both datasets. On the TenRec-QKVideo dataset, our model outperforms RL-based model (Pei et al., 2019) by 1.2 pp (3.919 v.s. 3.907) for 5-objective AUC sum. On the Kuaishou-ELive dataset, our model outperforms RL-based model by 0.55 pp (3.7522 v.s. 3.7467). These results verify our continuous gradient-based method is superior to RL policy gradient based methods. In a nutshell, our model consistently outperforms the best-performing baselines on both two datasets in a large margin in terms of different experimental settings.
Figure 3. The tradeoff curves between competing objectives of the different models on the Kuaishou Elive dataset.
5.5. Ablation study (RQ2)
We conduct comprehensive ablation study on Kuaishou-ELive dataset to investigate the impact of different AUC optimization losses and different model structures.
5.5.1. Different AUC optimization losses.
Other than our proposed differentiable ranking based AUC maximization methed, we compare several popular pairwise AUC surrogate losses i.e., pairwise logistic (Gao and Zhou, 2012), pairwise square (Gao et al., 2013) and the state-of-the-art instance-wise AUC maximization method AUCM (Yuan et al., 2021). Pairwise methods maximize AUC objective with logistic or square function surrogate to approximate pairwise $\mathbb{I}(f(\mathbf{x}_{+})\geq f(\mathbf{x}_{-}))$. For pairwise computation nature, these methods perform a $O(n^{2})$ time complexity during training process. Different from them, AUCM transforms pairwise maximization objective to an equivalent instance-wise min-max problem, which reduce the training time complexity to $O(n)$. Notwithstanding these efforts, the inherent inconsistency with the AUC metric persists. We compare the practical training speeds of different losses in Section A of the Appendix.
Different Losses AUC Sum Time-complexity
Pairwise Square 3.7415 $O(n^{2})$
Pairwise Logistic 3.7433 $O(n^{2})$
AUCM 3.7412 $O(n)$
Ours (Differentiable Ranking) 3.7522 $O(n\text{log}n)$
Table 2. Results on Kuaishou-ELive for different AUC-driven losses.
As shown in Tab. 2, our proposed differentiable ranking based method consistently outperforms other AUC losses. This is because our method eliminates the need of defining surrogate functions which makes it inevitably misaligned with AUC metric. This experiment fundamentally demonstrates the superiority of our method over existing methods.
5.5.2. Model components
As shown in Tab. 3, we conducted a comprehensive ablation study to evaluate the impact of different modules in our proposed method. We conduct three ablations in the relation aware module. First, when removing the self-attention-based alignment module (Ours w/o Self-Attention), we find the corresponding variant exhibits a 0.3pp drop in AUC sum (3.7494 v.s. 3.7522), indicating the indispensable role of relation modelling between different objective score encodings. Second, the variant without personalized feature (Ours w/o Personalized Feat.) exhibits a 0.6pp performance degradation (3.7464 v.s.3.7522), demonstrating their significant contribution in guiding score importance weighting. Third, in terms of the manner of introducing personalized information, in addition to cross-attention mechanism, we experiment with an alternative baseline that simply concatenating personalized features with self-attention outputs before linear projection (Ours w/o Cross-Attention). Our method outperforms this baseline by 0.1pp in AUC sum (3.7522 vs. 3.5712). We attribute this to linear projection’s inability to facilitate dense interactions between personalized features and score encodings which can be fulfilled by cross-attention’s dynamic weighting mechanism.
In relation-agnostic module, we conduct ablations on two sub-components inside it. First, we try to remove the gate mechanism (Ours w/o Gate Mechanism), which leads to substantial performance drop (1.6pp), confirming its critical function for denoising and importance score selection. In addition, the variant without linear fusion pathway Ours w/o Linear results in a 0.3pp AUC sum decrease (3.7493 v.s. 3.7522), suggesting that the first-order linear fusion can effectively enhance model robustness.
Method AUC Sum
[…]
Ours w/o Self-Attention 3.7494
Ours w/o Personalized Feat. 3.7464
Ours w/o Cross-Attention 3.7512
[…]
Ours w/o Gate Mechanism 3.7362
Ours w/o LinearPath 3.7493
Ours 3.7522
Table 3. Results on Kuaishou-Elive for different variants of our proposed structure with different components removed.
5.5.3. Pre-processing.
We also conduct experiments on different choices of number of buckets for score discretization. As shown in Tab. 4, we find the variant without discretization (No Disc.) results into a significant performance decrease (3.7522 $\rightarrow$ 3.7470), which demonstrates the necessity of discretization. We find different number of buckets present different performance, this may be because excessively fine-grained discretization approach the no-discretization situation, whereas overly coarse discretization make it lack discrimination between high and low scores. An appropriately number of buckets, i.e., 300 here, yields optimal performance.
#DiscretizationBuckets AUC Sum
No disc. 3.7470
100 3.7507
300 3.7522
600 3.7492
Table 4. Results on Kuaishou-Elive for different number of discretization buckets.
5.6. Online Results (RQ3)
Figure 4. Consistency analysis between the learned attention matrix $\mathbf{A}_{r}$ and ground-truth data distribution (label’s Spearman Rank Correlation $\rho$.
We deploy our method in the production environment of Kuaishou e-commerce live-streaming platform to conduct online A/B testing for 5 days. It is noteworthy that the online base model is MBCE. Compared with the base model. As shown in Tab. 5 our model achieves significant improvements across all objectives, e.g., 2.635% on core purchases metric and 0.451% on follow metrics, which verify the effectiveness of our proposed HarmonRank for offline-online consistent improvements.
Objectives Gain
Purchase +2.635%
Comment +3.034%
Follow +0.451%
Watch time +0.290%
Like +1.673%
Table 5. Online A/B Test Performance.
5.7. Analysis (RQ4)
5.7.1. Impact on Pareto frontier.
To validate whether our method can achieve Pareto improvement (i.e., improving at least one objective without deteriorating others) in multi-objective optimization, we compare the Pareto frontiers between our proposed method and Multi-objective BCE. Given that the loss weights assigned to different objectives significantly influence their corresponding AUC performance, we systematically adjusted the allocation of loss weights to obtain a series of AUC values under various trade-off scenarios. As seen in Fig. 3, our method consistently achieves superior trade-off curves over Multi-objective BCE (MBCE). Taking buy and follow objectives as examples, our method outperforms MBCE in terms of AUC for both objectives across all weight allocations. The similar phenomenon can be observed in trade-off curves between other objective pairs, including comment v.s. buy, comment v.s. follow and long view v.s. follow in Fig. 3.
5.7.2. Inter-objective alignment analysis.
To shed light on the factual degree of alignment between objectives (ground-truth), and which of them can be learned in inter-objective alignment modules (predicted), we visualize inter-objective correlations in two ways. For the ground-truth relations, we compute Spearman’s rank correlation coefficients $\mathbf{\rho}$ between different objectives over the labels, which measures the monotonic alignment between the ranks of two objective labels. For predicted relations, we visualize the attention weights inferred by our method’s relation module.
(22) $\mathbf{\rho}_{mn}$ $=\text{Spearmanr}(\mathbf{y}_{m},\mathbf{y}_{n})$
(23) $\mathbf{A}_{r}^{{mn}}$ $=\mathbf{A}_{r}[m,n]$
As shown in the Fig. 5, we present the $\rho_{buy,:}$ and $\mathbf{A}_{r}^{buy,:}$ between buy objective and all other objectives. It can be observed that the learned attention matrix exhibits strong alignment with label ranking, showing consistent patterns: comment ¿ long view ¿ follow ¿ like. This demonstrates that our inter-objective alignment module effectively captures latent dependencies. The particularly strong correlations of buy v.s. long view and buy v.s. comment are because the users with immediate purchase willingness tend to engage in prolonged viewing for product explanations and active questioning via comments.
Moreover, to quantitatively validate the consistency between learned patterns in the attention matrix and ground-truth inter-objective alignment, we try to compute the correlation strength between the ground-truth and predicted relations. To preserve more absolute correlation scales in attention matrix, we apply sigmoid instead of softmax normalization over the original dot product matrix. We compute the correlation between label Spearman’s $\mathbf{\rho}$ and predicted attention matrix $\mathbf{A}_{r}$.
(24) $r$ $=\text{Pearson}(\mathbf{\rho},\mathbf{A}_{r})$
Results in Fig .4 demonstrate remarkable consistency ($r=0.82$, $p=0.004$) between learned attentions and empirical alignment, confirming our method’s capability to accurately capture inter-objective alignment by our two-step ensemble paradigm.
Figure 5. Two views of the correlations between buy objective and other objectives: Spearman’s rank correlation ($\mathbf{\rho}$) of labels and learned attention matrix $\mathbf{A}_{r}$.
6. Conclusion
The proposed HarmonRank framework addresses two limitations in multi-objective ensemble problem in recommendation systems. For the lack of alignment to ranking, we formulate AUC computation as rank-sum problem and use differentiable ranking to enable end-to-end AUC optimization. For the lack of alignment between different objectives, we propose a two-step paradigm, first align then ensemble. This paradigm effectively align the shared parts between objectives. Extensive online and offline experimental results demonstrate significant improvements of our proposed HarmonRank over existing methods.
Appendix
In this appendix, we provide more experimental results of our proposed HarmonRank. Accordingly, we organize the appendix as follows.
  • In Section A, we present practical training speed our method and compared methods.
  • In Section B, we report the robustness of different methods under label skew situations.
Appendix A Practical training speed
As shown in Tab. 6, we compare the training speed among different AUC end-to-end losses. The training batch size is set to 10240 for all methods. The ranks of actual training speeds across different methods aligns with their theoretical computational complexity. Due to the pairwise nature of ‘PL’ (Pairwise Logistic), it exhibits the slowest training speed, while AUCM and MBCE methods achieve the fastest training owing to their instance-wise loss computation with $O(n)$ complexity. Although our method’s training speed lies between pairwise and instance-wise approaches, it delivers the best AUC performance across all methods.
Method Practical training speed Theoritical time complexity AUC Sum
MBCE 27.4K/s $O(n)$ 3.6682
PL 24.8K/s $O(n^{2})$ 3.7433
AUCM 26.9K/s $O(n)$ 3.7412
HarmonRank 26.0K/s $O(nlogn)$ 3.7522
Table 6. Comparison of practical training speed.
Appendix B Robustness
To validate the robustness of our method against label skew, we conducted extensive experiments on the Kuaishou E-live dataset. Among our five objectives, the buy objective exhibits the most severe label skew with a positive-to-negative ratio of $1:10^{3}$. To explore the performance of different methods under harder situations, we artificially strengthen the imbalance of buy objective by down-sampling positive samples to ratios of $1:10^{4}$ and $1:10^{5}$ respectively on training set only, remaining the test set unchanged.
As shown in the Tab. 7, when facing increasingly severe label skew, all methods exhibit expected AUC degradation. Nevertheless, our HarmonRank consistently maintains superior performance across all scenarios. Crucially, the performance drop of our method (-0.3%) is significantly smaller than those of MBCE and PL (9.8% and 4.3%), demonstrating our methods has exceptional robustness to label distribution skew.
We attribute this phenomenon to different degrees of dependable on negative samples of these methods. As positive samples become sparser, classifying negative samples becomes increasingly trivial, resulting in diminishing gradients from the overwhelming majority of easy negative examples. Consequently, methods like MBCE and PL that heavily rely on negative samples suffer significant performance degradation. In contrast, HarmonRank inherently rely less on negative samples for its rank-sum formulation, whose loss only computes ranks of positive samples and negative samples are utilized only during the differentiable ranking phase.
Methods Original ($1:10^{3}$) $10\times$ Skew ($1:10^{4}$) $100\times$ Skew ($1:10^{5}$)
[…]
Table 7. Robustness against label skew.